OpenStudy (anonymous):

Tell whether the sequence is arithmetic, geometric, or neither. Then supply the missing terms of the sequence.

OpenStudy (anonymous):

\[1,\frac{ 1 }{ 4 },\frac{ 1 }{ 9 }, \frac{ 1 }{ 16 }\] and two more missing terms.

OpenStudy (solomonzelman):

Actually.\(\Huge\color{blue}{ \bf \frac{1}{n^{2}} }\) first term \(\Huge\color{blue}{ \bf \frac{1}{1^{2}} }\) second term \(\Huge\color{blue}{ \bf \frac{1}{2^{2}} }\) third term \(\Huge\color{blue}{ \bf \frac{1}{3^{2}} }\)

OpenStudy (solomonzelman):

See?

OpenStudy (anonymous):

Thank you! just the first thing you wrote actually helped me figure this out... So the missing terms are 1/25, and 1/36

OpenStudy (solomonzelman):

Yes ! You are C♣rrect !